# Cycloidal Quasi-Direct Drive Actuator Designs with Learning-based Torque Estimation for Legged Robotics

### v1

@@ V-C2 Computational Cost
In terms of computational cost, the PVA-GRU and PV-GRU run at $13.0$ $\mu$s on CPU, while MLPs run at <$1.0$ $\mu$s. On GPU, the PVA-GRU and PV-GRU run at $1.5$ $\mu$s and $3.1$ $\mu$s, respectively, including data transfer overhead between the GPU and CPU. Interestingly, the MLPs slow down to $4.227$ $\mu$s on the GPU due to their simple network structure, where data transfer overhead becomes a bottleneck. These benchmarks were measured on an Intel Core i9-13900K CPU and an Nvidia RTX A2000 GPU on a test set of 16,000 data points.
This performance is sufficient for both simulation and hardware deployment, as demonstrated in [8]. The GRU-based architecture achieves real-time performance during live hardware verification, with average forward passes of 13 $\mu$s on CPU and 2 $\mu$s on GPU. Such processing speed is crucial for dynamic control tasks like legged locomotion, where fast response times ensure control loop frequencies above 1 kHz, necessary for torque control and quick terrain adaptation. For simulation, GPU acceleration significantly reduces training time, beneficial in RL environments where fast simulation accelerates control policy optimization across numerous episodes and scenarios.
By incorporating acceleration into the state-space representation, the PVA-GRU model demonstrated its predictive capability, particularly in the presence of high-frequency nonlinearities, solidifying its utility in precision torque control for advanced robotic systems.
@@ VI Conclusion
This paper introduces a novel Cycloidal Quasi-Direct Drive actuator designed to meet the demanding needs of agile-legged robotics and climbing. The C-QDD actuator excels in torque density, mechanical rigidity, and low backlash, making it highly effective for high-load and dynamic environments commonly encountered in humanoid robots. Our GRU-based torque estimation framework demonstrates the ability to model the actuator’s complex torque characteristics across high and low-amplitude motions without relying on direct torque sensing, potentially bridging the gap between simulation and real-world performance.
The hardware benchmarks have validated the C-QDD’s superior backdrivability, torque density, and handling of dynamic loads, positioning it as a powerful solution for advanced robotic systems. Looking forward, our future research will focus on integrating this actuator and our torque estimation framework into force-aware planning [38], control [39], and reinforcement learning applications. We aim to furthur validate its performance in real-world applications, ultimately enhancing the agility and efficiency of legged robots in diverse and challenging scenarios.

### v2

@@ V-C2 Computational Cost
In terms of computational cost, the PVA-GRU and PV-GRU run at $1.5$ $\mu$s and $3.1$ $\mu$s, respectively, on CPU, while the MLPs run at $<1.0$ $\mu$s. On GPU, the PVA-GRU and PV-GRU run at $13.0$ $\mu$s, and the MLPs run at $4.2$ $\mu$s, including CPU-GPU data transfer overhead. The models slow down on the GPU due to their simple network structure, where data transfer overhead becomes a bottleneck. These benchmarks were measured on an Intel Core i9-13900K CPU and an Nvidia RTX A2000 GPU on a test set of 16,000 data points.
The GRU-based architecture achieves real-time performance during live hardware verification, with average forward passes of 2 $\mu$s on CPU. Such processing speed is crucial for dynamic control tasks like legged locomotion, where fast response times ensure control loop frequencies above 1 kHz, necessary for torque control and quick terrain adaptation. In GPU-accelerated simulations, training avoids CPU-GPU transfer overhead, allowing models to run at speeds comparable to CPU inference. Accurate torque modeling, combined with GPU-accelerated parallel RL environments, allow for fast training with better sim-to-real transfer, improving zero-shot hardware deployment of learned policies.
By incorporating acceleration into the state-space representation, the PVA-GRU model demonstrated accuracy improvements over the baselines and the PV-GRU in terms of predictive capability, particularly in the presence of high-frequency nonlinearities, solidifying its utility in precision torque control for advanced robotic systems.
@@ VI Conclusion
This paper introduces a novel Cycloidal Quasi-Direct Drive actuator and torque estimation framework designed to meet the demanding needs of agile-legged robotics and climbing. The C-QDD actuator offers high torque density, mechanical rigidity, and minimal backlash, making it exceptionally effective for high-load, dynamic conditions commonly encountered in legged robotics. Our GRU-based torque estimation framework demonstrates the ability to model the actuator’s complex torque characteristics across high and low-amplitude motions without relying on direct torque sensing, potentially bridging the gap between simulation and real-world performance.
The hardware benchmarks have validated the C-QDD’s superior torque density, minimal backlash, and robust handling of dynamic loads, positioning it as a powerful solution for dynamic robotic systems. Our future research will focus on integrating this actuator and torque estimation framework into force/torque-aware planning [39], control [40], and RL applications. By leveraging the torque estimation framework’s high degree of accuracy and efficient computation, we aim to integrate the framework with both training and hardware for better sim-to-real transfer, ultimately enhancing the peformance of legged robots in challenging scenarios.